OpenStudy (imreallyasiren):

Can someone please help me out on this question? I've seen multiple people answer this before but they all don't seem correct.

OpenStudy (imreallyasiren):

Given the function f(x) = 4(2)^x, Section A is from x = 1 to x = 2 and Section B is from x = 3 to x = 4. Part A: Find the average rate of change of each section. Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other.

OpenStudy (imreallyasiren):

Here is the question

OpenStudy (legomyego180):

\[f(x)=4(2)^x\] Plug in. \[f(1)=4(2)^x\] \[f(2)=4(2)^x\] \[f(3)=4(2)^x\] \[f(4)=4(2)^x\] Rate of change = Slope = Rise over run = Y-value / x-value

OpenStudy (imreallyasiren):

okay so in order it would be 8, 16, 32, and 64?

OpenStudy (legomyego180):

Remember y value over x-value the first would be 8/1 second 16/2 third 32/3 fourth 64/4 dont forget to simplify

OpenStudy (imreallyasiren):

okay so for section A i got 8 for both and for b i got 32

OpenStudy (imreallyasiren):

okay i think i figured it out thank you